# Optimistic Estimation of Convergence in Markov Chains with the Average-Mixing Time

### v2

@@ 4.3   Infinite graphs with controlled growth
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph [^6] With the convention that a disconnected graph has infinite diameter., $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
Condition 4 (Connection graph with polynomial growth).
@@
Let $p,q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
$\mathcal{J}_{p,\xi}\leq G^{1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q/p}\left\|\pi\right\|^{1-1/p}_{1-1/p}.$
@@
$\begin{split}\mathcal{J}_{p}^{(s)}&=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{X}}P^{s}(x,x^{\prime})^{1-1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\sum_{x^{\prime}\in\mathcal{S}_{x}^{(s)}}\left(\frac{1}{\mathcal{S}_{x}^{(s)}}\right)^{1-1/p}\\ &=\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}{\mathcal{S}_{x}^{(s)}}^{1/p}\leq\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}(s^{q}G)^{1/p},\\ \end{split}$
where the first inequality stems from the uniform distribution maximizing the $\ell_{1-1/p}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
Observe that
$\lim_{p\to\infty}\left\|\pi\right\|_{1-1/p}^{1-1/p}=\lim_{p\to\infty}\left[t^{\sharp}_{\mathsf{mix}}(\xi)\right]^{q/p}=1,$
that is, for large values of $p$, the upper bound of Lemma 4.5 still approaches unity.
Example 4.1 (Linear growth and stationary distribution with polynomially decaying tail).
We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\eta\in(1/8,1/4)$ such that $\pi(x)\leq\frac{C}{x^{1+\eta}}$ for any $x\in\mathcal{X}$. Taking $p=2(1+1/\eta)$, it holds that
$\sum_{x\in\mathcal{X}}\pi(x)^{1-1/p}\leq C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{(1+\eta)(1-1/p)}}=C^{1-1/p}\sum_{x\in\mathcal{X}}\frac{1}{x^{1+\eta/2}}=C^{2\frac{\eta+2}{\eta+1}}\zeta(1+\eta/2),$
where $\zeta$ is the Riemann Zeta function. When the connection graph of $P$ enjoys linear growth, that is $q=1$, an application of Theorem 3.4, yields that for $\delta\in(0,1/2)$, a trajectory length of
$n\geq\widetilde{\mathcal{O}}\left(\frac{t_{\mathsf{mix}}^{11/10}}{\delta^{2}}\right),$
is sufficient to obtain $\widehat{t_{\mathsf{mix}}}^{\sharp}(2\delta/5)\geq t^{\sharp}_{\mathsf{mix}}(\delta/2)$ with probability at least $1-\delta$.
4.4   General finite space
When the state space $\mathcal{X}$ is finite, the graph-related methods discussed in the previous section remain applicable. Alternatively, a more direct approach yields the following bound,
$\sup_{s\in\mathbb{N}}\mathcal{J}_{p}^{(s)}\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1-1/p}=\left|\mathcal{X}\right|^{2/p},$ (10)
agnostic of the mixing properties of $P$. Note that (10) matches the heuristic lower bound $\left\|\pi\right\|_{1-1/p}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
@@ 4.4   General finite space
$n\geq 1+\frac{2^{8}et_{\mathsf{mix}}}{\xi^{2}\varepsilon^{2}}\max\left\{\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+2\log_{2}\left|\mathcal{X}\right|\right),9\left(\left\lceil\log_{2}\frac{4}{3\xi}\right\rceil+1\right)\log\left(\frac{4t_{\mathsf{mix}}}{\delta}\left\lceil\log_{2}\frac{1}{\xi}\right\rceil\right)\right\},$
@@
Proof.
We plug (10) in Theorem 3.4 and optimize for $p$. The function $p\mapsto p\left|\mathcal{X}\right|^{2/p}$ is minimized for $p=2\log\left|\mathcal{X}\right|$, whence the corollary. ∎
@@ 4.5   The two-point space
$\begin{split}P_{p,q}^{t}&=1^{\intercal}\pi_{p,q}+\frac{(1-p-q)^{t}}{p+q}\begin{pmatrix}p&-p\\ -q&q\end{pmatrix},\\ \end{split}$
@@
$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\qquad d_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
@@
$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{(1+\eta)^{2}}{2\eta}\doteq r_{\eta}(t),$
@@
5   Proofs
@@
The function $x\mapsto x/\log x$ is increasing on $[e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
@@
5.1   Proof of Lemma 2.3
The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1993, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959). For simplicity, we assume that $n=2Bs$, where $B,s\in\mathbb{N}$. For $b\in[B]$, we write
$\begin{split}X^{[2b]}&=X^{[2b]}_{1},\dots,X^{[2b]}_{s}=X_{(2b-1)s+1},\dots,X_{2bs},\\ X^{[2b-1]}&=X^{[2b-1]}_{1},\dots,X^{[2b-1]}_{s}=X_{(2b-2)s+1},\dots,X_{(2b-1)s},\\ \end{split}$
@@
Figure 2: Decomposing the process into blocks.
@@
it holds that $22(B-1)\beta(s)\leq\delta/$ 2. As a consequence, the deviation probability at (11) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
@@
and invoking Lemma 5.1, it is enough that
@@
$\begin{split}d^{\sharp}(t)&\doteq\sum_{x\in\mathcal{X}}\pi(x)\left\|\delta_{x}P^{t}-\pi\right\|_{\mathsf{TV}}\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\left|P^{t}(x,x^{\prime})-\pi(x^{\prime})\right|\\ &=\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\\ &\stackrel{{\scriptstyle(\spadesuit)}}{{\leq}}\frac{1}{2}\sum_{x,x^{\prime}\in\mathcal{X}}\pi(x)\pi(x^{\prime})\frac{(1-\gamma_{\star})^{t}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\sum_{x,x^{\prime}\in\mathcal{X}}{\sqrt{\pi(x)\pi(x^{\prime})}}\\ &=\frac{1}{2}(1-\gamma_{\star})^{t}\left\|\pi\right\|_{1/2}.\\ \end{split}$
@@
$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|\lambda_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
@@
5.3   Proof of Theorem 3.1
For $s\in\mathbb{N}$,
@@
We further decompose the argument in the second expectation as
$\begin{split}&\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|\\ &\leq\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}\right|+\left|\frac{N^{(s)}_{x}\pi(x^{\prime})}{\lfloor(n-1)/s\rfloor}-\pi(x)\pi(x^{\prime})\right|\\ &=2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\\ \end{split}$
Noting that by stationarity,
@@
it follows from Jensen’s inequality that
@@
We proceed to bound the two variance terms. We first treat $\operatorname{Var}_{\pi}\left[N_{x}\right]$ for a fixed $x\in\mathcal{X}$. This will in turn yield bounds for general $\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]$ and $\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]$. The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (see e.g. Rio (1999, (1.6))) as,
From Rio (1993, Theorem 1) [^7] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant. (see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)), for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
where $\alpha$ is Rosenblatt’s strong mixing coefficient [^8] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999). (Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (see e.g. (Bradley, 2005, (1.11))). Plugging in (13), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right),\end{split}$
and for a general $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}^{(s)}\right]&\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\right)\\ &\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor\pi(x)^{1-1/p}\mathcal{B}^{(s)}_{p}.\\ \end{split}$
We verify (see e.g. Wolfer and Kontorovich (2021, Lemma 6.1)) that the process consisting of a sliding window over pairs,
@@
forms a Markov chain over the state space =$\mathcal{E}=\left\{x,x^{\prime}\in\mathcal{X}\colon P(x,x^{\prime})>0\right\}$, called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2024) of the original chain, with transition operator
$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),$
@@
$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2}\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2},\end{split}$
and it follows from the Markov property that for $t\in\mathbb{N},t<n-1$,
@@
More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s,\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
@@
As a consequence
$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}\right),$
where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)^{s}P(x,x^{\prime})$. We further bound
$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-2}\beta(s(t-1))^{1/p}\\ &\leq\frac{1}{2}\left(1-\left\|\pi\right\|_{2}^{2}\right)+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st))^{1/p}\\ &\leq\frac{3}{2}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-3}\beta(st)^{1/p}\leq\frac{3}{2}\mathcal{B}_{p}^{(s)},\end{split}$
@@
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\sqrt{\left\|Q^{(s)}\right\|_{1-1/p}^{1-1/p}\frac{3}{2}\mathcal{B}_{p}^{(s)}}+2\sqrt{\left\|\pi\right\|_{1-1/p}^{1-1/p}\mathcal{B}_{p}^{(s)}}\right),\end{split}$
which yields the first inequality in the theorem. The second inequality can be obtained from observing that $\left\|Q^{(s)}\right\|_{1-1/p}\geq\left\|\pi\right\|_{1-1/p}$. ∎
5.4   Proof of Corollary 3.1
@@
$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$

### v4

@@ 4.2  Infinite graphs with controlled growth
where $\rho\colon\mathcal{X}\times\mathcal{X}^{2}\to\mathbb{N}\cup\left\{\infty\right\}$ denotes the diameter of a directed graph, which is the maximum eccentricity of any vertex in the graph with the convention that a disconnected graph has infinite diameter, $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ is the sub-graph of $(\mathcal{X},\mathcal{E})$ induced by the vertices in $\mathcal{S}$, and the supremum is taken over all $\mathcal{S}$ such that $(\mathcal{X},\mathcal{E})|_{\mathcal{S}}$ has diameter at most $s$.
Condition 1 (Connection graph with polynomial growth).
@@
Example 4.1 (Birth and death chains).
It is immediate from its definition that the connection graph associated with the Chebyshev type random walk developed in Example 2.1 has bounded growth with $G=2$ and $q=1$.
@@
Let $p\in\mathbb{R}_{+}\cup\{+\infty\},q\in[1,\infty),\xi\in(0,1)$ and let $P\in\mathcal{W}(\mathcal{X})$ be a $\pi$-stationary transition operator. If the connection graph of $P$ has polynomial growth of degree $q$ with factor $G$, it holds that
$\mathcal{J}_{p,\xi}\leq G^{1+1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q(1+1/p)}\left\|\pi\right\|^{1-1/p}_{(1-1/p)/2}.$
@@
[…]
where the first inequality stems from the uniform distribution maximizing the $\ell_{(1-1/p)/2}$ quasi-norm, and the second inequality follows from the polynomial growth assumption. ∎
Example 4.2 (Graph with polynomial growth and stationary distribution with polynomially decaying tail).
We let $\mathcal{X}=\mathbb{N}$, and suppose that there exists $C\in\mathbb{R}_{+}$, and $\alpha\in(1,\infty)$ such that $\pi(x)\leq\frac{C}{x^{1+\alpha}}$ for any $x\in\mathcal{X}$. It holds that
$\sum_{x\in\mathcal{X}}\pi(x)^{(1-1/p)/2}\leq C^{(1-1/p)/2}\zeta((1+\alpha)(1-1/p)/2),$
where $\zeta$ is the Riemann Zeta function.
  1. 1.
    In the uniformly ergodic setting, Lemma 4.1 specializes to
    $\mathcal{J}_{\infty,\xi}\leq GCt^{\sharp}_{\mathsf{mix}}(\xi)^{q}\zeta((1+\alpha)/2)^{2},$
    and treating $C,G,\alpha$ as constants, an upper bound on the sample complexity for a $(1/4,\delta)$-PAC bound on $t^{\sharp}_{\mathsf{mix}}(\xi)$ is given by
    $n\geq\widetilde{\mathcal{O}}\left(t_{\mathsf{mix}}^{q+1}/\xi^{2}\right).$
  2. 2.
    In the non-uniformly setting, we may instead optimize the following product
    $\min_{p\geq 1}\left\{\mathcal{B}_{p}G^{1+1/p}t^{\sharp}_{\mathsf{mix}}(\xi)^{q(1+1/p)}\left\|\pi\right\|^{1-1/p}_{(1-1/p)/2}\right\}.$
4.3  Finite space
When the state space $\mathcal{X}$ is finite and $P$ is supported on $(\mathcal{X},\mathcal{E})$, the graph-related methods discussed in the previous section remain applicable. Alternatively, a cruder—graph agnostic—approach yields the following bound,
$\sqrt{\mathcal{J}_{\infty}^{(s)}}\leq\sum_{(x,x^{\prime})\in\mathcal{E}}\left(\frac{1}{\left|\mathcal{X}\right|^{2}}\right)^{1/2}=\left|\mathcal{X}\right|.$
Note that (4.3) matches the heuristic lower bound $\left\|\pi\right\|_{(1-1/p)/2}^{2(1-1/p)}$ in the worst-case where $P$ is bistochastic.
@@ 4.3  Finite space
$n\geq 1+4t_{\mathsf{mix}}\frac{1+\left\lceil\log_{2}\frac{1}{\xi(1-\varepsilon)}\right\rceil}{\xi^{2}\varepsilon^{2}}\max\left\{C_{\mathsf{u.erg.}}^{2}\left|\mathcal{X}\right|^{2},576\log\frac{4t_{\mathsf{mix}}\left\lceil\log_{2}1/\xi\right\rceil}{\delta}\right\},$
@@
Comparison with related work
Since the criteria for convergence are different, the problems of estimating the relaxation or worst-case mixing time of a chain to multiplicative accuracy and the problem at hand are not strictly speaking comparable, but it is instructive to inspect the sample complexities. The problem of estimating the relaxation time of a Markov chain to multiplicative error has sample complexity of the order of $\widetilde{\Theta}(t_{\mathsf{mix}}/\pi_{\star})$ (Hsu et al., 2019; Wolfer and Kontorovich, 2024), while a known upper bound on the complexity of estimating the mixing time is of the order of $\widetilde{\mathcal{O}}(\left|\mathcal{X}\right|t_{\mathsf{mix}}/\pi_{\star})$ (Wolfer, 2024). Crucially, the above-mentioned problem all suffer a dependency on the minimum stationary probability $\pi_{\star}$. When $\pi_{\star}\propto\exp(-\left|\mathcal{X}\right|)$, the problem becomes intractable. Corollary 4.1 yields the superior upper bound $\widetilde{\mathcal{O}}(t_{\mathsf{mix}}\left|\mathcal{X}\right|^{2}/\xi^{2})$, agnostic of $\pi_{\star}$ and enjoying an at most linear dependency on the number of parameters of the transition operator.
@@ 4.4  The two-point space
[…]
@@
$\begin{split}d_{p,q}^{\sharp}(t)&=2\frac{pq}{(p+q)^{2}}\left|1-p-q\right|^{t},\hskip 18.49988ptd_{p,q}(t)=\frac{p\lor q}{p+q}\left|1-p-q\right|^{t}.\end{split}$
@@
$\frac{d_{p,q}(t)}{d^{\sharp}_{p,q}(t)}=\frac{1+\eta}{2\eta}\doteq r_{\eta}(t),$
@@
5  Conclusion
To assess convergence in Markov chains from a single trajectory of observations, recent literature has focused on estimating the worst-case mixing time or the spectral gaps associated with the transition operator. These mixing parameters lead to concentration inequalities that are almost as powerful as those in the iid setting. However, their existence requires strong assumptions on the chain such as uniform or geometric ergodicity, and they are notoriously hard to estimate from the data, with known lower bounds inversely proportional to the minimum stationary probability $\pi_{\star}$. As a result, the estimation problem is only statistically tractable for moderately small space sizes and when the tail of the stationary distribution does not decay too rapidly. In this paper, we relaxed these necessary conditions and limitations by considering the average-mixing time instead. Over finite spaces, our analysis showed that the average-mixing time is easier to estimate than its worst-case counterpart, and we could obtain an upper bound on the sample complexity of the problem that is independent of the rate of decay of the stationary distribution. Moreover, in contradistinction to worst-case mixing time estimation, our analysis extends to countable state spaces and sub-geometrically $\beta$-mixing Markov chains under natural structural assumptions. The trade-off for accessing an easier convergence estimation rate is a logarithmic degradation in the available concentration inequalities.
Appendix A Extension to non-stationary Markov chains
Although our results are generally stated assuming a stationary start, it is possible to generalize them to non-stationary chains as follows. Let $\mu\in\mathcal{P}(\mathcal{X})$ absolutely continuous with respect to $\pi$. Inspired by the proof of Paulin (2015, Proposition 3.10), an application of Hölder’s inequality with $q,r\geq 1$ with $1/q+1/r=1$ yields,
Appendix B Connection with $\beta$-mixing
$\lim_{s\to\infty}\beta(s)=0.$
In particular, a stationary countable-state Markov chain $X$ is $\beta$-mixing if and only if it is irreducible and aperiodic (Bradley, 2005, Theorem 3.2).
Lemma B.1 (Davydov (1974, Proposition 1)).
$\beta(s)=\sup_{t\in\mathbb{N}}\sum_{x\in\mathcal{X}}\mu^{(t)}(x)\left\|e_{x}P^{s}-\mu^{(s+t)}\right\|_{\mathsf{TV}}.\\$
[…]
thus the average-mixing time is essentially the “stationary $\beta$-mixing time” of the chain [^2] Similarly, the worst-case mixing time is directly connected to the notion of $\phi$-mixing (Davydov, 1974). . Note that with the convention $P^{0}=I$, we obtain
$\beta(0)=1-\left\|\pi\right\|_{2}^{2},$
which depends solely on the stationary properties of the chain.
Appendix C Proofs
@@
The function $x\mapsto x/\log x$ is increasing on $e,\infty)$ and in this range, it holds that $x/\log x\geq e>2$. Hence, for $B\geq 2A\log A$, it holds that
@@
C.1  Proof of Lemma 2.1
The proof is standard, and based on the blocking technique credited to Bernstein and further developed in the series of papers (Yu, 1994; Eberlein, 1984; Volkonskii and Rozanov, 1959). For simplicity, we assume that $n=2Bs$, where $B,s\in\mathbb{N}$. For $b\in[B]$, we write
[…]
@@
Figure 4: Decomposing the process into Bernstein blocks.
@@
it holds that $2(B-1)\beta(s)\leq\delta/2$. As a consequence, the deviation probability at (14) is bounded by $\delta$ when the trajectory length $n$ satisfies the inequality,
@@
and invoking Lemma C.1, it is enough that
@@
[…]
@@
$\left|\frac{P^{t}(x,x^{\prime})}{\pi(x^{\prime})}-1\right|\leq\left|1-\gamma_{\star}\right|^{t}\sqrt{\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x)\sum_{i=2}^{\left|\mathcal{X}\right|}f_{i}^{2}(x^{\prime})}.$
@@
C.3  Proof of Theorem 3.1
The proof of Theorem 3.1 will rely on the following lemma.
Lemma C.2.
Let $X_{1},\dots,X_{n}$ be a Markov chain over a countable state space $\mathcal{X}$ with stationary distribution $\pi$. For any $x\in\mathcal{X}$, and any $p\in[1,\infty)$, it holds that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right)\leq 4(n-1)\pi(x)^{1-1/p}\mathcal{B}_{p},\end{split}$
where $N_{x}=\sum_{t=1}^{n-1}\delta\left[X_{t}=x\right]$ counts the number of visits to state $x\in\mathcal{X}$.
Proof.
The sequence $(\delta\left[X_{t}=x\right])_{t\in\mathbb{N}}$ being stationary at second order, we expand the variance of the sum (Rio, 1999, (1.6)) as,
From Rio (1993, Theorem 1) [^3] This generalization of Ibragimov’s covariance inequality (Ibragimov, 1962) also appears in Davydov (1968) but with a weaker constant.—see also Rio (1999, (1.12b)) and Doukhan (2012, Theorem 3)— for a Hölder triplet $p,r,q\geq 1,\frac{1}{p}+\frac{1}{q}+\frac{1}{r}=1$ and any $t\in\mathbb{N}\cup\left\{0\right\}$,
where $\alpha$ is Rosenblatt’s strong mixing coefficient [^4] We note that $\alpha$ is defined as twice Rosenblatt’s coefficient in Rio (1999). (Rosenblatt, 1956), we wrote $\left\|X\right\|_{\pi,r}=\left(\mathbb{E}_{\pi}\left|X\right|^{r}\right)^{1/r}$ for a real random variable $X$, and for the second inequality, we relied on the inequality $2\alpha(t)\leq\beta(t)$ (Bradley, 2005, (1.11)). Plugging in (15), it follows that
$\begin{split}\operatorname{Var}_{\pi}\left[N_{x}\right]&\leq 2(n-1)\beta(0)^{1/p}\pi(x)^{1-1/p}+4\sum_{t=1}^{n-2}(n-1-t)\beta(t)^{1/p}\pi(x)^{1-1/p}\\ &\leq 4(n-1)\pi(x)^{1-1/p}\left(\frac{\beta(0)^{1/p}}{2}+\sum_{t=1}^{n-2}\beta(t)^{1/p}\right).\end{split}$
Moving on to the proof of Theorem 3.1, for $s\in\mathbb{N}$, by definition of $\widehat{\beta}(s)$ in (8),
@@
With a symmetry argument, we further decompose the argument in the second expectation as
$\begin{split}\sum_{x,x^{\prime}\in\mathcal{X}}\left|\frac{N^{(s)}_{x}N^{(s)}_{x^{\prime}}}{\lfloor(n-1)/s\rfloor^{2}}-\pi(x)\pi(x^{\prime})\right|&\leq 2\sum_{x\in\mathcal{X}}\left|\frac{N^{(s)}_{x}}{\lfloor(n-1)/s\rfloor}-\pi(x)\right|.\end{split}$
By stationarity,
@@
and it follows from Jensen’s inequality that
@@
We proceed to bound the two variance terms. From Lemma C.2, for a fixed $x\in\mathcal{X}$ and any $s<n$, noting that the mixing coefficient $\beta^{(s)}(t)$ pertaining to the $s$-skipped chain satisfies $\beta^{(s)}(t)=\beta(st)$, we obtain
[…]
The process consisting of a sliding window over pairs,
@@
forms a Markov chain (Wolfer and Kontorovich, 2021, Lemma 6.1) over the state space
$\mathcal{E}=\left\{(x,x^{\prime})\in\mathcal{X}^{2}\colon P(x,x^{\prime})>0\right\},$
called the Hudson expansion (Kemeny and Snell, 1983; Wolfer and Watanabe, 2024) of the original chain, with transition operator
$\overline{P}(e=(x_{1},x_{2}),e^{\prime}=(x_{1}^{\prime},x_{2}^{\prime}))=\delta\left[x_{2}=x_{1}^{\prime}\right]P(x_{2},x_{2}^{\prime}),\forall(e,e^{\prime})\in\mathcal{E}^{2},$
@@
$\begin{split}\overline{\beta}(0)&=1-\left\|\overline{\pi}\right\|_{2}^{2},\\ \overline{\beta}(1)&=\beta(0)=1-\left\|\pi\right\|_{2}^{2}.\end{split}$
Moreover, it follows from the Markov property that for $t\in\mathbb{N}$,
@@
More generally, for $s\in\mathbb{N},t\in\mathbb{N}\cup\left\{0\right\}$, $1\leq s\leq n-1$, and $0\leq t<\lfloor(n-1)/s\rfloor$,
@@
An application of Lemma C.2 on the Hudson expanded chain yields that
$\operatorname{Var}_{\pi}\left[N_{xx^{\prime}}^{(s)}\right]\leq 4\left\lfloor\frac{n-1}{s}\right\rfloor Q^{(s)}(x,x^{\prime})^{1-1/p}\left(\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\overline{\beta}^{(s)}(t)^{1/p}\right),$
where we recall that $Q^{(s)}(x,x^{\prime})=\pi(x)P^{s}(x,x^{\prime})$. We further bound
$\begin{split}\frac{\overline{\beta}^{(s)}(0)^{1/p}}{2}+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\overline{\beta}^{(s)}(t)^{1/p}&=\frac{1}{2}\left(1-\left\|Q^{(s)}\right\|_{2}^{2}\right)+\sum_{t=1}^{\lfloor(n-1)/s\rfloor-1}\beta(s(t-1))^{1/p}\\ &\leq 1/2+\sum_{t=0}^{\lfloor(n-1)/s\rfloor-2}\beta(st)^{1/p}\leq 1/2+\mathcal{B}_{p}^{(s)},\end{split}$
@@
$\begin{split}\mathbb{E}_{\pi}\left|\widehat{\beta}(s)-\beta(s)\right|&\leq\sqrt{\frac{1}{\lfloor(n-1)/s\rfloor}}\left(\left\|Q^{(s)}\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{1/2+\mathcal{B}_{p}^{(s)}}+2\left\|\pi\right\|_{(1-1/p)/2}^{(1-1/p)/2}\sqrt{\mathcal{B}_{p}^{(s)}}\right).\end{split}$
Finally, the claim can be obtained from observing that $\left\|Q^{(s)}\right\|_{(1-1/p)/2}\geq\left\|\pi\right\|_{(1-1/p)/2}$. ∎
C.4  Proof of Lemma 3.1
@@
$\begin{split}\mathcal{B}_{p}^{(s)}=\beta_{0}^{1/p}\sum_{t=0}^{\lfloor(n-1)/s\rfloor-1}\exp\left(-\beta_{1}st/p\right)&\leq\frac{\beta_{0}^{1/p}}{1-\exp(-\beta_{1}s/p)}.\end{split}$